The wage rate of labor is Rs. 6 and price of capital is Rs. 2. The marginal product of labor is 16 while marginal product of capital is 4. Can a firm be operating at equilibrium?       


1
Expert's answer
2021-08-17T17:46:11-0400

Solution:

A firm will operate at equilibrium where: "\frac{MP_{L} }{P_{L} } = \frac{MP_{K} }{P_{K} }"

 

The equilibrium point is the point at which the firm is producing the maximum amount of output at a given cost.

MPL = 16

PL = 6

MPK = 4

PK = 2

At equilibrium: "\frac{16 }{6 } = \frac{4 }{2 }"


At equilibrium: 2.67 = 2


"\frac{MP_{L} }{P_{L} } > \frac{MP_{K} }{P_{K} }"

The firm is not operating at equilibrium since "\frac{MP_{L} }{P_{L} } > \frac{MP_{K} }{P_{K} }", that is they are not equal, which shows that the marginal benefit of additional labor exceeds the marginal cost of capital.